How many 5-card hands having exactly 3 aces and 2 other cards can be dealt?
Answer by stanbon(75887)

How many 5-card hands having exactly 3 aces and 2 other cards can be dealt?
-------------
# of ways to pick 3 aces: 4C3 = 4
# of ways to pick 2 other cards: 48C2 = 48*47/1*2 = 1128
----------------
Total # of 5-card hands with 3 aces and 2 other cards: 4*1128 = 4512
===================
Cheers,
Stan H.